The value of planck's constant is 6.63×10^-34 js
The velocity of the light is 3×10^8 ms^-1
Which Value is closest to the wavelength in nanometers of a quantum of light with frequency of 8×10^15s^-1

Grade:11

1 Answers

Arun
5 years ago
For light, the energy E is given by E = hf and also by E = hc/λ, where h is Planck's constant, f is the frequency, c is the speed of light, and λ is the wavelength. So, we have: 
hf = hc/λ 
f = c/λ 
fλ = c 
λ = c/f 
λ = (3.0 x 10^8) / (6 x 10^15) 
λ = 5 x 10^-8 m 
λ = 5 x 10^1 nm